OpenStudy (anonymous):

determine the zeros of the function and their multiplicity....y=x to the 5th power(x-5)to the 2nd power(x-6)to the 4th power

OpenStudy (anonymous):

Do you mean this: \[\sqrt{x+7} = x - 13\] ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So firstly square both the sides... \[(\sqrt{x + 7})^2 = (x- 13)^2\] \[\implies x + 7 = (x - 13)^2\] Can you go forward now??

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

Just do the calculations and post your answer here.. In case of any difficulty, let me know..